During an experiment, a signal from a spaceship reached the ground station in five minutes. What was the distance of the spaceship from the ground station? The signal travels at the speed of light, that is, 3 × 108 m s-1.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Speed= 3 × 108 m/s
Time= 5 min = 5 *60 = 300 seconds.
Distance= Speed * Time
Distance= 3 * 108m/s * 300 secs. = 9 x 10^10m

Thanks